OpenStudy (anonymous):

I know you aren't suppose to give direct answers but I need a quick answer. What is the value of the y variable in the solution to the following system of equations? 3x + y = 14 –3x + 2y = 1 A)5 B)3 C) y can be any number as there are infinitely many solutions to the system. D)There is no y value as there is no solution to the system.

OpenStudy (whpalmer4):

Add the two equations together and you'll get an equation in terms of y. That can be easily solved.

OpenStudy (anonymous):

if you just need answers quick, try this http://www.wolframalpha.com/input/?i=+3x+%2B+y+%3D+14+%2C%E2%80%933x+%2B+2y+%3D+1

OpenStudy (anonymous):

Thank you. I know it is easy but i'm sorta in a situation and I need to answer all of them like asap.

OpenStudy (whpalmer4):

Add down the columns, you get 3x-3x + y + 2y = 14+1 3y = 15 y=5 was that so hard?

OpenStudy (anonymous):

No.

OpenStudy (whpalmer4):

Using WolframAlpha is handy, no argument there, but if you're doing homework or a test, it becomes more a measure of whether you can type in the problems and copy out the answers rather than whether you know the material. Believe me, this is material you'll use quite frequently before you graduate, and skipping past it when you don't know it will be a long-term detriment. Anyhow, good luck with the time pressure!

OpenStudy (anonymous):

With similar linear systems of equations, just remember you need to cancel one of the variables. For example, if the first equation was x+y = 14, you would've had to multiply 3(x) + 3(y) = 3(14).....and THEN add down the columns to solve for Y.